A ladder 42 feet long is place so that it will reach a window 30 feet high (first building) on one side of a street; if it is turned over, its foot being held in position, it will reach a window 2o5 feet high (second building) on the other side of the street. How wide is the street from the building to building?
Answer #1
I assume that the second window is 25 feet high.

Draw a diagram. Let the foot of the ladder be x feet from the first building, and y feet from the second building. Using the good old Pythagorean Theorem, we have

x^2 + 30^2 = 42^2
y^2 + 25^2 = 42^2

so, x=29.39 and y=33.75, so the street is roughly 63 feet wide.
answered by: ---------------------------------
Answer #2
l
i l
i l
i l
i l
i____l
answered by: lucely
Answer #3
Thank you!! I trust you. :D
I hope this is right. :D

*This is my illustration

l
i l
i l
i l
i l
i____l


-the 'l' is the 1st building.
-the 'i' is the 2nd building.
-the "backslash()" is the ladder.
-the "underscore(_)" is the street.

>is this right? :D
answered by: ******
Answer #4
Your diagram is wrong. The ladder does not reach from one building to the other. Its foot is somewhere in the middle of teh street, and it pivots so that it touches the first building at a height of 30', and the other at a height of 25'. Read the problem.
answered by: K. C
Answer #5
woah! it doesn't look the way i made it. :( it's just look like letter 'N'. if it's fix.